Pls Help! : Each interior angle of a regular n-gon is 156. Calculate the value of n. ( using the sum of the interior angles of an n-gon)
Answer by stanbon(48535)

Each interior angle of a regular n-gon is 156. Calculate the value of n. ( using the sum of the interior angles of an n-gon)
-----------------------------
EQUATION:
(Sum of the interior angles)/(number of sides) = measure of one interior angle
------------------------------
[(n-2)*180]/n = 156
180n - 360 ] = 156n
24n = 360
n = 15
There are 15 sides.
======================
Cheers,
Stan H.